OpenStudy (anonymous):

Linear Algebra problem. Let A be an 2x2 matrix over the real numbers. 1. The subset {A^2, A^5, A^11} is always linearly dependant. 2. The subset {I, A, A^2} is always linearly dependant. 3. It is possible that the subset {A^2, A^5, A^11} is linearly independant. 4. It is possible that the subset {I, A^2, A^5, A^11} is linearly independant. 5. It is possible that the subset {I, A, A^2} is linearly independant. Appreciate your help! notice I means the identity matrix.

OpenStudy (goformit100):

@ash2326

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Coolsector - It is pretty hard question, huh?

OpenStudy (anonymous):

no you are asked the opposite

OpenStudy (anonymous):

linearly dependant.

OpenStudy (anonymous):

yes lol sorry

OpenStudy (anonymous):

It is really hard to find such an example. I'm trying....

OpenStudy (anonymous):

Theorticlly It should be possible somehow because you don't exceed the 4 element limit (the dimension of the 2x2 matrices is 4). If you provide one answer it will provide an answer almsot to all the others.

OpenStudy (anonymous):

are you here? still trying?

OpenStudy (anonymous):

@imron07 do you have an idea?

OpenStudy (anonymous):

It seems like this post is not going to be resolved.

OpenStudy (experimentx):

try asking it here http://math.stackexchange.com/

OpenStudy (experimentx):

read FAQ first http://math.stackexchange.com/faq

OpenStudy (experimentx):

if you get response ... also let me know.

OpenStudy (anonymous):

thank you @experimentX

OpenStudy (experimentx):

honestly i would also like to know it.

OpenStudy (anonymous):

It means that (2) is right. But I still want to understand why... I can't row-reduce x^2....and such terms.

OpenStudy (experimentx):

try putting matrices as {{1,2},{3,4}}

OpenStudy (anonymous):

I did..... I tried anything... like 20 times

OpenStudy (anonymous):

Nothing works. all are depended

OpenStudy (experimentx):

aI + bA + cA^2 = 0 ... aI + A(bI + cA) = 0 ------------------- try playing around!!

OpenStudy (anonymous):

Okay @experimentX, I have the answers and I promised to include them here as I find them so here it is. By the "Cayley-Hamilton" theorem we can actually say that every square matrix over a commutative ring satisfies its own characteristic equation. That is, \[A^n = \lambda_1(I) + \lambda_2(A)\] Which means the the dimension of the subspace spanned by the A powers has at most dimension 2. It follows that every set with 3 or more members is linearly dependent. Lets see some concrete example: see (!) and (2) above. These must be linearly depenedent subsets because every one of them is a subset of the A powers. As I mentioned earlier the basis of a subset spanned by the A powers has at most dimension 2. I'll show you why. Lets look at the second one. We have the subset \[{I, A, A^2}\] Now the identity matrix is simply A^ 0. so we have a subset of A powers. I will show you now that A^2 can be expressed with only I and A. If I show that It's true, we are done because this is exactly the defintion of linearly dependent subset. A^2 is a sqaure matrix so we can use "Cayley-Hamilton" theorem. characteristic polynomial is given by \[p_A(λ)=λ^2−(a+d)λ+(ad−bc)\]so the Cayley–Hamilton theorem states that \[p(A)=A^2−(a+d)A+(ad−bc)I\] and therefore : \[A^2=(a+d)A-(ad−bc)I\] and we are done! The same thing can be done with any subset of A powers. I hope everything is clear. Let me know if you have questions.

OpenStudy (anonymous):

Also @Coolsector , you might find this one interesting.

OpenStudy (anonymous):

ty